OpenStudy (coconutjj):

(1/9)^x=3*sqrt(27) Solve for x, without log

OpenStudy (anonymous):

Well, prime factorize both, I guess

OpenStudy (anonymous):

\[ \frac{1}{9}=3^{-2} \]And\[ 3\sqrt{27} = 3^{1+\frac32} \]

OpenStudy (anonymous):

So \[ -2x=\frac 52 \]I guess

OpenStudy (solomonzelman):

\(\large\color{black}{ (1/9)^x=3\sqrt{27} }\) \(\large\color{black}{ (9)^{-x}=9\sqrt{3} }\) \(\large\color{black}{ (9)^{-x-1}=\sqrt{3} }\) \(\large\color{black}{ (9)^{-x-1}=9^{1/4}}\)

OpenStudy (solomonzelman):

making sense?

OpenStudy (solomonzelman):

I am getting the same result, I think.